OpenStudy (anonymous):

Can someone help me solve the second equation for x^2 and use substitution for the two values of y in the problem: x^2+y^2-16y+39=0 y^2-x^2-9=0

OpenStudy (whpalmer4):

as you have x^2 + y^2 - 16y + 39 = 0 for the first equation (terms in x^2, y^2, and y) and y^2 - x^2 -9 = 0 for the second (terms in x^2 and y^2) I would inclined to solve the second for x^2, and substitute that into the first one.

OpenStudy (whpalmer4):

after you do that, it's pretty smooth sailing to the solution. note that there are 3 solutions, not 2...

OpenStudy (anonymous):

I'm really bad at setting these up, could you by chance help me? @whpalmer4

OpenStudy (whpalmer4):

y^2 - x^2 -9 = 0 if we add x^2 to both sides, we get y^2 - x^2 + x^2 -9 = 0 + x^2 x^2 = y^2 - 9 right?

OpenStudy (whpalmer4):

now we rewrite the other equation, putting (y^2 - 9) wherever we find x^2 x^2 + y^2 - 16y + 39 = 0 (y^2 - 9) + y^2 - 16y + 39 = 0 see how you do solving that for y and report back :-)

OpenStudy (anonymous):

I get y=3,5! @whpalmer4

OpenStudy (anonymous):

add the two equations together, the x^2 term will drop out

OpenStudy (anonymous):

2y^2 - 16y + 30 = 0 solve for y

OpenStudy (anonymous):

or if you prefer y^2 - 8y + 15=0 factors as (y-3)(y-5) = 0 so y = 3 or y = 5

OpenStudy (anonymous):

Ah, thank you both so much! Is that how you solve the equation for x^2, though? Because it says "Solve the equation for x^2" and then says "Use substitution to solve for two values of y"

OpenStudy (whpalmer4):

that's true, those are the correct values of y, but if you plug them into the equation and solve for x, there are actually 3 values of x that satisfy the equation. I believe there would be 4, except that one of the values of x is 0, and -0 is indistinguishable from 0 :-)

OpenStudy (whpalmer4):

we did solve the second equation for x^2: it equals y^2 - 9.

OpenStudy (whpalmer4):

then we substituted y^2 - 9 for x^2 in the first equation and found the values of y that satisfied the equation.

OpenStudy (whpalmer4):

Here are some plots of the intersection/solutions:

OpenStudy (anonymous):

When I plug in for the x values, I get 4i, -4i, would that be right?

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (whpalmer4):

no, made a mistake somewhere.

OpenStudy (whpalmer4):

y^2 - x^2 - 9 = 0 let's use y = 3 (3)^2 - x^2 - 9 - 0 9 - x^2 - 9 = 0 -x^2 = 0 x = 0 (that's the one that gives us an odd number of solutions) you try y = 5...

OpenStudy (anonymous):

I get x= +/- 4!

OpenStudy (whpalmer4):

yes, those are the correct answers!

OpenStudy (anonymous):

Thanks so much for all your help!